# Learning Fused State Representations for Control from Multi-View Observations

### v1

@@ 7 Discussion
The proposed MFSC framework advances MVRL by tackling key challenges in learning task-relevant representations from multi-view observations. By integrating bisimulation metric learning with a masking-based latent reconstruction auxiliary task, MFSC enhances robustness and adaptability in environments with redundant, distracting, or incomplete inputs. These improvements are particularly beneficial for applications such as autonomous robotics, self-driving systems, and intelligent surveillance, where robust perception under real-world uncertainties is critical.
From an ethical standpoint, biases in sensor placement, data modalities, or patterns of missing observations could introduce unintended disparities in learned policies. Future research should explore strategies to mitigate these biases, ensuring consistent and reliable deployment across diverse scenarios. Moreover, as reinforcement learning gains traction in safety-critical domains, rigorous validation and improved interpretability of learned policies are essential to prevent unintended consequences and enhance trustworthiness.
@@ Appendix A Proofs
$(1-c)|V^{*}(\vec{o})-V^{*}(\phi(\vec{o}))|\leq g(\vec{o},\tilde{d})+\frac{\gamma}{1-\gamma}\max_{u\in\mathcal{O}^{1}\times\mathcal{O}^{2}\times\cdots\times\mathcal{O}^{k}}g(u,\tilde{d}),$ (14)
@@
As $\varepsilon\rightarrow 0$, the optimal value function of the aggregated MDP converges to the original value function. By defining a learning error for $\phi$, $\mathcal{L}:=\sup_{\vec{o}_{i},\vec{o}_{j}\in\mathcal{O}^{1}\times\mathcal{O}^{2}\times\cdots\times\mathcal{O}^{k}}\big{|}||\phi(\vec{o}_{i})-\phi(\vec{o}_{j})||_{1}-{\tilde{d}}(\vec{o}_{i},\vec{o}_{j})\big{|}$, we can also update the bound in Lemma 4.2 to incorporate $\mathcal{L}:|V^{*}(\vec{o})-V^{*}(\phi(\vec{o}))|\leq\frac{2\varepsilon+2\mathcal{L}}{(1-\gamma)(1-c)}$.
@@ Appendix B Additional Related Work Discussion
In this section, we provide a more detailed discussion on state representation learning in RL and a detailed comparison between our method and other related works.
@@ C.1 Hyperparameters
Cube spatial size $12\times 12$ $12\times 12$
@@
Critic target update frequency $2$ $2$
Actor update frequency $2$ $2$
@@
Temperature learning rate $1\times 10^{-4}$ $1\times 10^{-4}$
@@
Cube spatial size $12\times 12$ $12\times 12$
@@ E.1 Results on DeepMind Control Suite
Cartpole Swingup 832$\pm$29 717$\pm$134 758$\pm$155 865$\pm$15 864$\pm$2 862$\pm$9 --
Cheetah Run 679$\pm$33 19$\pm$14 502$\pm$120 710$\pm$52 915$\pm$6 841$\pm$87 537$\pm$60
Finger Spin 977$\pm$7 289$\pm$104 873$\pm$145 860$\pm$104 978$\pm$10 594$\pm$136 985$\pm$2
Walker walk 909$\pm$40 41$\pm$17 818$\pm$261 770$\pm$371 968$\pm$3 959$\pm$27 940$\pm$8
@@ E.3 Visualization and Analysis on Noisy View
To investigate whether bisimulation can assist MFSC in countering the interference of noise, we conducted ablation experiments under the setting of interference views. The results confirmed that bisimulation effectively aids MFSC in mitigating the impact of noise, allowing it to maintain its performance even in the presence of disruptive noise. This demonstrates the robustness of MFSC when enhanced with bisimulation, highlighting its ability to focus on relevant features while disregarding irrelevant noise.
Figure 13: Ablation study on robust view.
E.4 Hyperparameter Sensitivity Analysis
@@
E.5 Results on Different Network Architectures
@@
E.6 Comparison of training resources
@@
View Concatenating Vanilla PPO 3.8h 2.67ms 850MB --
RAD 3.8h 2.69ms 758MB --
@@ Appendix F Algorithm
feature_a_norm = nn.functional.normalize(feature_a, dim=1)
feature_b_norm = nn.functional.normalize(feature_b, dim=1)

### v4

@@ Acknowledgements
The proposed MFSC framework advances MVRL by tackling key challenges in learning task-relevant representations from multi-view observations. By combining bisimulation-based representation learning with a masking-based latent reconstruction auxiliary task, MFSC improves robustness and adaptability in environments characterized by redundant, distracting, or incomplete inputs. These improvements are particularly beneficial for applications such as autonomous robotics, self-driving systems and other scenarios where reliable perception is essential.
Ethical considerations include potential biases from sensor placement, modality imbalance, or observation dropout, which may affect policy fairness and generalizability. Future work should explore bias mitigation strategies and improve interpretability, particularly in safety-critical deployments.
@@ Appendix A Proofs
[…]
@@
[…]
@@ Appendix B Additional Related Work Discussion
In this section, we provide a more detailed discussion on state representation learning in RL and a detailed comparison between our method and other related work.
@@ C.1 Hyperparameters
Cube spatial size $\times 12 12×12$$\times 12$
@@
Critic target update frequency 22 22
Actor update frequency 22 22
@@
Temperature learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
@@
Cube spatial size $\times 12 12×12$$\times 12$
@@ E.1 Results on DeepMind Control Suite
[…]
@@ E.3 Visualization and Analysis on Noisy View
To investigate whether bisimulation can assist MFSC in countering the interference of noise, we conducted ablation experiments under the setting of interference views. The results of Fig.LABEL:robust_view_ablation confirm that bisimulation effectively aids MFSC in mitigating the impact of noise, allowing it to maintain its performance even in the presence of disruptive noise. This demonstrates the robustness of MFSC when enhanced with bisimulation, highlighting its ability to focus on relevant features while disregarding irrelevant noise.
(a) (a) Ablation study on robust view. (b) Ablation study on varying view counts.
E.4 Ablation Study on Varying View Counts
We have added an ablation study that varies the number of views, and the results are presented in FigLABEL:fig:view_num. It shows that using any single view out of the three performs worse than the full multi-view setup. In particular, View 3 — a top-down perspective — suffers from severe occlusion, resulting in the lowest performance among the three. This highlights the limitations of relying on individual views and underscores the effectiveness of our approach in extracting and fusing complementary information across views. Overall, these results validate the necessity and benefit of multi-view learning in achieving more robust and informative state representations.
E.5 Quantitative Analysis on Bisimulation Metric
To provide a more quantitative analysis of the bisimulation metric, we include the training curve of the bisimulation loss and additionally measure the mutual information $I\left(z_{L}^{0};s\right)$ between the final fused embeddings $z_{L}^{0}$ and the ground-truth states $s$ using the MINE [^4] https://github.com/gtegner/mine-pytorch method. As illustrated in Fig.14, with the progression of training and the convergence of bisimulation loss, we observe a consistent increase in mutual information. This indicates that the model not only optimizes the bisimulation criterion but also gradually constructs a task-relevant representation space aligned with the true environment states.
Figure 14: Quantitative analysis on bisimulation metric.
E.6 Hyperparameter Sensitivity Analysis
@@
E.7 Results on Different Network Architectures
@@
E.8 Results of the Comparison with Additional MVRL Algorithm
We compare our method with the model-based MVRL algorithm MOSER (Wan et al., 2024). MOSER seeks the optimal viewpoint for learning task representations under multiple views. Its implementation code is available at this link [^5] https://github.com/yixiaoshenghua/MOSER. The corresponding experimental results are illustrated in Fig.18. Our experimental results show that our approach outperforms the MOSER algorithm in both tasks, demonstrating the effectiveness of our method.
Figure 18: Performance comparison of an MB-MVRL algorithm MOSER.
E.9 Comparison of Training Resources
@@
View Concatenating Vanilla PPO 3.8h 2.67ms 850MB -
RAD 3.8h 2.69ms 758MB -
@@ Appendix F Algorithm
feature_a_norm = $n$.functional.normalize(feature_a, dim=1)
feature_b_norm = $n$.functional.normalize(feature_b, dim=1)